OpenStudy (anonymous):

Someone help me pleaseeee Solve the equation 4p^2+16p=0

OpenStudy (emmamink):

p = 0

OpenStudy (solomonzelman):

\(\large\color{black}{ 4p^2+16p=0 }\) \(\large\color{black}{ 4p(p+4)=0 }\)

OpenStudy (solomonzelman):

So either 4p=0, or (p+4)=0. your answers are?

OpenStudy (anonymous):

The answer is 4 and 0 I just didn't know how to get that answer